# Parallel transport in rotating frames and projective holonomic quantum computation

### v2

@@ II.D   Central rotations as gauge symmetries
A rotational transformation that belongs to the center of the group of unitary operators on $\mathcal{H}$ is a phase factor multiple of the identity operator,
@@
Such a transformation does not affect the system’s dynamics because it only adds a global phase to state vectors and thus leave states and observables unchanged. However, it adds a multiple of the identity to the Hamiltonian,
@@
The parallel transport condition (3) is not invariant under such gauge transformations. This is evident from the fact that $H_{t}$ and $H_{t}+\epsilon_{t}\mathbbm{1}$ cannot both satisfy the parallel transport condition in the laboratory frame, despite generating identical dynamics. In the next section, we will derive a condition for parallel transport that remains invariant under the symmetry transformations described by Eq. (17).
@@ III.A   Projective gates
on the gates of $\mathcal{R}$. The projective gates form a group that acts on density operators with support in $\mathcal{R}$ as
@@
where $\Gamma$ is any representative of $\bar{\Gamma}$. Abstractly, the group of projective gates on $\mathcal{R}$ is the holomorphic isometry group on the projective space formed by the unit rank density operators with support in $\mathcal{R}$.
@@ III.B   The projective Stiefel-Grassmann bundle
The projective Stiefel manifold is the manifold of orbits under the action by the center on the Stiefel manifold,
@@
Thus, the projective Stiefel manifold consists of equivalence classes of $n$-frames in $\mathcal{H}$, where two frames belong to the same class if and only if they differ by a phase factor,
@@
The projective Stiefel-Grassmann bundle is a principal fiber bundle with symmetry group the projective unitary group $\mathrm{PU}(n)=\mathrm{U}(n)/\mathrm{Z}(n)$. The projective unitary group is a Lie group with Lie algebra $\mathfrak{pu}(n)=\mathfrak{u}(n)/\mathfrak{z}(n)$. [^8] The Lie algebra $\mathfrak{pu}(n)$ is canonically isomorphic to the Lie algebra of trace-free skew-Hermitian $n\times n$ matrices $\mathfrak{su}(n)$, although $\mathrm{PU}(n)$ is not isomorphic to the special unitary group $\mathrm{SU}(n)$. The canonical isomorphism is $\mathfrak{pu}(n)\ni A+\mathfrak{z}(n)\mapsto A-\operatorname{tr}(A)/n\in\mathfrak{su}(n)$, with inverse $\mathfrak{su}(n)\ni A\mapsto A+\mathfrak{z}(n)\in\mathfrak{pu}(n)$.
There is a unique $\mathfrak{pu}(n)$-valued connection $\mathcal{PA}$ on the projective Stiefel manifold making the diagram
@@
commutative for every $n$-frame $V$. The vertical arrows are the tangent maps of the projections onto the respective orbit spaces. Every tangent vector $Y$ at $\bar{V}$ is the image of a tangent vector $X$ at $V$, and $Y$ is projectively horizontal, that is, $\mathcal{P}\mathcal{A}(Y)=0$, if and only if $\mathcal{A}(X)\in\mathfrak{z}(n)$.
@@ III.C   Projective holonomic gates and projectively parallel transporting Hamiltonians
Let $\mathcal{R}_{t}$ be a closed curve in $\mathcal{G}(n;\mathcal{H})$ at $\mathcal{R}$, and let $\bar{V}$ be any projective frame for $\mathcal{R}$. There exists a unique curve of projective $n$-frames $\bar{V}_{t}$ over $\mathcal{R}_{t}$ that emanates from $\bar{V}$ and whose velocity belongs to the kernel bundle of $\mathcal{PA}$ at all times [21]. This curve is the projective horizontal lift of $\mathcal{R}_{t}$ that starts at $\bar{V}$.
Suppose $V$ is an $n$-frame representing $\bar{V}$. Then $\bar{V}_{t}$ can be lifted to a curve of frames $V_{t}$ for $\mathcal{R}_{t}$ starting at $V$. Any other lift of $\bar{V}_{t}$ to $\mathcal{V}(n;\mathcal{H})$ takes the form $e^{i\theta_{t}}V_{t}$ for some smooth family of phases $\theta_{t}$. That $\bar{V}_{t}$ is projectively horizontal is equivalent to $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$. If we write $|v_{j;t}\rangle$ for the $j$th component of $V_{t}$, we can equivalently formulate the projective horizontality condition as
@@
for some real numbers $\epsilon_{t}$. Note that $V_{t}$ is conventionally horizontal when $\epsilon_{t}=0$, which we can always achieve by choosing $\theta_{t}$ appropriately. We define the projective holonomy of $\mathcal{R}_{t}$ as the projective unitary operator of $\mathcal{R}$ represented by $V_{\tau}V^{\dagger}$. This definition is independent of both the choice of projective horizontal lift $\bar{V}_{t}$ of $\mathcal{R}_{t}$ and the choice of lift $V_{t}$ of $\bar{V}_{t}$ to $\mathcal{V}(n;\mathcal{H})$. The projective holonomy transforms states with support in $\mathcal{R}$ in the same way as the conventional holonomy.
A Hamiltonian $H_{t}$ with time-translation operator $U_{t}$ is said to be projectively parallel translating if for some (and hence any) $n$-frame $V$ for $\mathcal{R}$, the curve $V_{t}=U_{t}(V)$ satisfies $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$ and accordingly represents a projective horizontal curve. Writing $|v_{j}\rangle$ for the $j$th component of $V$, and $|v_{j;t}\rangle$ for $U_{t}|v_{j}\rangle$, the projective parallel transport condition becomes
@@
for some real-valued $\epsilon_{t}$; c.f. [9, Eq. (25)]. Unlike the conventional parallel transport condition (3), the projective parallel transport condition is gauge invariant in the sense that if $H_{t}$ is projectively parallel transporting, so is $H_{t}+\epsilon_{t}\mathbbm{1}$ for every real-valued function $\epsilon_{t}$.
@@ III.D   Projective parallel transport in a rotating frame picture
Suppose $V_{t}$ is a curve of $n$-frames. Let $V_{t}^{\textsc{rf}}=R_{t}V_{t}$ be its rotated counterpart, and $A_{t}=i\dot{R}_{t}R_{t}^{\dagger}$ be the rotational potential. Then
$V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)\iff V_{t}^{{\textsc{rf}}\dagger}(d_{t}+iA_{t})V_{t}^{{\textsc{rf}}}\in\mathfrak{z}(n).$ (27)
@@
$\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}|v_{l;t}^{\textsc{rf}}\rangle=i\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n),$ (28)
@@
$\langle v_{k;t}^{\textsc{rf}}|H_{t}^{\textsc{rf}}-A_{t}|v_{l;t}^{\textsc{rf}}\rangle=\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n).$ (29)
@@ III.E   The isoholonomic inequality for projective holonomic gates
We define the isoholonomic bound of the projective gate $\bar{\Gamma}$ as the minimum of the isoholonomic bounds of the gates that represent $\bar{\Gamma}$,
@@
Now, suppose $\mathcal{R}_{t}$ is a loop in $\mathcal{G}(n;\mathcal{H})$ with projective holonomy $\bar{\Gamma}$. By the isoholonomic inequality for gates [20],
@@
$L(\bar{\Gamma})=\min_{0\leq k\leq n}\left\{\sqrt{\sum_{j=1}^{n}|\theta_{j}-\theta_{k}|\big{(}2\pi-|\theta_{j}-\theta_{k}|\big{)}}\,\right\}.$ (32)
To see this, first observe that
$L(\bar{\Gamma})=\min_{0\leq\theta<2\pi}\{L(e^{-i\theta}\Gamma)\}.$ (33)
@@
$L(e^{-i\theta}\Gamma)^{2}=\sum_{j=1}^{n}|\theta_{j}-\theta|\big{(}2\pi-|\theta_{j}-\theta|\big{)}.$ (34)
@@ III.F   An execution time estimate for projective holonomic quantum gates
$\tau[H_{t};\Gamma]\geq\tau[H_{t};\bar{\Gamma}]=\frac{L(\bar{\Gamma})}{\langle\!\langle\sqrt{I(H_{t};\mathcal{R}_{t})}\,\rangle\!\rangle}.$ (36)
@@
In Ref. [20] we demonstrated that if $\mathcal{R}$ has codimension at least $n$, every gate $\Gamma$ can be generated by a parallel transporting Hamiltonian $H_{t}$ in the time $\tau[H_{t};\Gamma]$. [^9] Whether one can reduce the codimension even more is an open question. Since this holds for any representative $\Gamma$ of $\bar{\Gamma}$ that satisfies $L(\Gamma)=L(\bar{\Gamma})$, we can conclude that if $\mathcal{R}$ has codimension at least $n$, every projective gate $\bar{\Gamma}$ can be generated by a projective parallel transporting Hamiltonian $H_{t}$ in the time $\tau[H_{t};\bar{\Gamma}]$. [^10] This is not as obvious as it may seem because $H_{t}$ and all its shifted versions $H_{t}+\epsilon_{t}\mathbbm{1}$ evolve $\mathcal{R}$ along the same path and with the same speed. Thus, if $H_{t}$ generates $\bar{\Gamma}$ suboptimally, that is, in a time longer than $\tau[H_{t};\bar{\Gamma}]$, the same is true for $H_{t}+\epsilon_{t}\mathbbm{1}$.
IV   Summary

### v3

@@ II.D   Central rotations as gauge symmetries
A rotation transformation belonging to the center of the group of unitary operators on $\mathcal{H}$ is a phase factor multiple of the identity operator,
@@
Such a transformation leaves states and observables unchanged and thus does not affect the dynamics of the system. However, it does add a multiple of the identity to the Hamiltonian:
@@
The parallel transport condition (3) is not invariant under such gauge transformations. This is evident from the fact that $H_{t}$ and $H_{t}+\epsilon_{t}\mathbbm{1}$ cannot both satisfy the parallel transport condition in the laboratory frame (unless $\epsilon_{t}=0$), even though they generate identical dynamics. In the next section we will derive a parallel transport condition that remains invariant under the symmetry transformations described by Eq. (17).
@@ III.A   Projective gates
on the gates of $\mathcal{R}$. The projective gates form a group which acts on density operators with support in $\mathcal{R}$ as
@@
where $\Gamma$ is an arbitrary representative of $\bar{\Gamma}$. Abstractly, the group of projective gates on $\mathcal{R}$ is the holomorphic isometry group on the projective space formed by the unit rank density operators with support in $\mathcal{R}$.
@@ III.B   The projective Stiefel-Grassmann bundle
The projective Stiefel manifold is the manifold of orbits under the action of the center on the Stiefel manifold,
@@
Thus, the projective Stiefel manifold consists of equivalence classes of $n$-frames in $\mathcal{H}$, where two frames belong to the same class if and only if they differ by a phase factor:
@@
The projective Stiefel-Grassmann bundle is a principal fiber bundle with symmetry group the projective unitary group $\mathrm{PU}(n)=\mathrm{U}(n)/\mathrm{Z}(n)$. The projective unitary group is a Lie group with Lie algebra $\mathfrak{pu}(n)=\mathfrak{u}(n)/\mathfrak{z}(n)$. [^8] The Lie algebra $\mathfrak{pu}(n)$ is canonically isomorphic to the Lie algebra of trace-free skew-Hermitian $n\times n$ matrices $\mathfrak{su}(n)$, although $\mathrm{PU}(n)$ is not isomorphic to the special unitary group $\mathrm{SU}(n)$. The canonical isomorphism is $\mathfrak{pu}(n)\ni A+\mathfrak{z}(n)\mapsto A-\operatorname{tr}(A)/n\in\mathfrak{su}(n)$, with the inverse $\mathfrak{su}(n)\ni A\mapsto A+\mathfrak{z}(n)\in\mathfrak{pu}(n)$.
There is a unique $\mathfrak{pu}(n)$-valued connection $\mathcal{PA}$ on the projective Stiefel manifold that makes the diagram
@@
commutative for every $n$-frame $V$. The vertical arrows are the tangent maps of the projections onto the respective orbit spaces. Each tangent vector $Y$ at $\bar{V}$ is the image of a tangent vector $X$ at $V$, and $Y$ is projectively horizontal, that is, $\mathcal{P}\mathcal{A}(Y)=0$, if and only if $\mathcal{A}(X)\in\mathfrak{z}(n)$.
@@ III.C   Projective holonomic gates and projectively parallel transporting Hamiltonians
Assume $\mathcal{R}_{t}$ is a closed curve in $\mathcal{G}(n;\mathcal{H})$ at $\mathcal{R}$, and let $\bar{V}$ be any projective frame for $\mathcal{R}$. There exists a unique curve of projective $n$-frames $\bar{V}_{t}$ over $\mathcal{R}_{t}$ which starts at $\bar{V}$ and whose velocity vector belongs to the kernel bundle of $\mathcal{PA}$ at all times [23]. This curve is the projective horizontal lift of $\mathcal{R}_{t}$ starting at $\bar{V}$.
Suppose $V$ is an $n$-frame representing $\bar{V}$. Then $\bar{V}_{t}$ can be lifted to a curve of frames $V_{t}$ for $\mathcal{R}_{t}$ starting at $V$. Any other lift of $\bar{V}_{t}$ to $\mathcal{V}(n;\mathcal{H})$ takes the form $e^{i\theta_{t}}V_{t}$ for some smooth family of phases $\theta_{t}$. That $\bar{V}_{t}$ is projectively horizontal is equivalent to $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$. If we write $|v_{k;t}\rangle$ for the $k$th component of $V_{t}$, we can equivalently formulate the projective horizontality condition as
@@
for some real numbers $\epsilon_{t}$. Note that $V_{t}$ is conventionally horizontal when $\epsilon_{t}=0$, which we can always achieve by choosing $\theta_{t}$ appropriately. We define the projective holonomy of $\mathcal{R}_{t}$ as the projective unitary operator of $\mathcal{R}$ represented by $V_{\tau}V^{\dagger}$. This definition is independent of both the choice of the projective horizontal lift $\bar{V}_{t}$ of $\mathcal{R}_{t}$ and the choice of the lift $V_{t}$ of $\bar{V}_{t}$ to $\mathcal{V}(n;\mathcal{H})$. The projective holonomy transforms states with support in $\mathcal{R}$ in the same way as the conventional holonomy.
A Hamiltonian $H_{t}$ with time translation operator $U_{t}$ is said to be projectively parallel translating if, for some (and hence any) $n$-frame $V$ for $\mathcal{R}$, the curve $V_{t}=U_{t}(V)$ satisfies $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$ and thus represents a projective horizontal curve. Writing $|v_{k}\rangle$ for the $k$th component of $V$ and $|v_{k;t}\rangle$ for $U_{t}|v_{k}\rangle$, the projective parallel transport condition becomes
@@
for some real-valued $\epsilon_{t}$; c.f. [9, Eq. (25)]. Unlike the conventional parallel transport condition (3), the projective parallel transport condition is gauge invariant in the sense that if $H_{t}$ is projectively parallel transporting, so is $H_{t}+\epsilon_{t}\mathbbm{1}$ for any real-valued function $\epsilon_{t}$.
@@ III.D   Projective parallel transport in a rotating frame picture
Assume $V_{t}$ is a curve of $n$-frames. Let $V_{t}^{\textsc{rf}}=R_{t}V_{t}$ be its rotated counterpart, and let $A_{t}^{\textsc{rf}}=i\dot{R}_{t}R_{t}^{\dagger}$ be the rotational potential. Then
$V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)\iff V_{t}^{{\textsc{rf}}\dagger}(d_{t}+iA_{t}^{\textsc{rf}})V_{t}^{{\textsc{rf}}}\in\mathfrak{z}(n).$ (27)
@@
$\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}^{\textsc{rf}}|v_{l;t}^{\textsc{rf}}\rangle=i\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n),$ (28)
@@
$\langle v_{k;t}^{\textsc{rf}}|H_{t}^{\textsc{rf}}-A_{t}^{\textsc{rf}}|v_{l;t}^{\textsc{rf}}\rangle=\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n).$ (29)
@@ III.E   The isoholonomic inequality for projective holonomic gates
We define the isoholonomic bound of the projective gate $\bar{\Gamma}$ as the minimum of the isoholonomic bounds of the gates representing $\bar{\Gamma}$,
@@
Now, suppose $\mathcal{R}_{t}$ is a loop in $\mathcal{G}(n;\mathcal{H})$ with projective holonomy $\bar{\Gamma}$. By the isoholonomic inequality for gates,
@@
$L(\bar{\Gamma})=\min_{0\leq k\leq n}\left\{\sqrt{\sum_{l=1}^{n}|\theta_{l}-\theta_{k}|\big{(}2\pi-|\theta_{l}-\theta_{k}|\big{)}}\,\right\}.$ (32)
To see this, first note that
$L(\bar{\Gamma})=\min_{0\leq\theta<2\pi}\big{\{}L(e^{-i\theta}\Gamma)\big{\}}.$ (33)
@@
$L(e^{-i\theta}\Gamma)^{2}=\sum_{l=1}^{n}|\theta_{l}-\theta|\big{(}2\pi-|\theta_{l}-\theta|\big{)}.$ (34)
@@ III.F   An execution time estimate for projective holonomic quantum gates
$\tau_{\textsc{qsl}}[H_{t};\Gamma]\geq\tau_{\textsc{qsl}}[H_{t};\bar{\Gamma}]=\frac{L(\bar{\Gamma})}{\langle\!\langle\sqrt{I(H_{t};\mathcal{R}_{t})}\,\rangle\!\rangle}.$ (36)
@@
Reference [22] showed that if $\mathcal{R}$ has codimension at least $n$, then any gate $\Gamma$ can be generated by a parallel transporting Hamiltonian $H_{t}$ in the time $\tau_{\textsc{qsl}}[H_{t};\Gamma]$. Since this holds for any representative $\Gamma$ of $\bar{\Gamma}$ that satisfies $L(\Gamma)=L(\bar{\Gamma})$, we can conclude that if $\mathcal{R}$ has codimension at least $n$, then any projective gate $\bar{\Gamma}$ can be generated by a projective parallel transporting Hamiltonian $H_{t}$ in the time $\tau_{\textsc{qsl}}[H_{t};\bar{\Gamma}]$. [^9] This is not as obvious as it might seem, because $H_{t}$ and all its shifted versions $H_{t}+\epsilon_{t}\mathbbm{1}$ evolve $\mathcal{R}$ along the same path and at the same rate. So if $H_{t}$ generates $\bar{\Gamma}$ suboptimally, that is, in a time longer than $\tau_{\textsc{qsl}}[H_{t};\bar{\Gamma}]$, the same is true for $H_{t}+\epsilon_{t}\mathbbm{1}$.
IV   Tight implementations
Whether the evolution time estimate (10) can be saturated using a parallel transporting Hamiltonian is highly relevant for the design of time-optimal implementations of holonomic gates. We call an implementation that saturates this estimate a tight implementation. Here, we revisit the proof in Ref. [22] that if the codimension of the computational space is at least as large as the dimension, then any quantum gate can be tightly implemented using a parallel transporting Hamiltonian.
Let $\Gamma$ be any gate. The parallel transporting Hamiltonian constructed in [22] to implement $\Gamma$ is of the form
$H=\sum_{k=1}^{n}H_{k;t}.$ (37)
The summands $H_{k;t}$ are effective qubit Hamiltonians acting on pairwise perpendicular $2$-dimensional subspaces $\mathcal{E}_{k}$, each spanned by a normalized eigenvector $|v_{k}\rangle$ of $\Gamma$,
$\Gamma|v_{k}\rangle=e^{i\theta_{k}}|v_{k}\rangle,$ (38)
and a vector perpendicular to the computational space. [^10] Thus $\mathcal{H}$ is orthogonally decomposed as $\mathcal{H}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}\oplus\cdots\oplus\mathcal{E}_{n}\oplus\mathcal{F}$. The subspace $\mathcal{E}_{k}$ is spanned by an eigenvector of $\Gamma$ and a vector in the orthogonal complement of $\mathcal{R}$. The summand $H_{k;t}$ leaves $\mathcal{F}$ and all the $\mathcal{E}_{l}$’s stationary except $\mathcal{E}_{k}$. As before, we assume that $0\leq\theta_{k}<2\pi$. A key result shown in [22] is that $H_{k;t}$ can be chosen such that
  • (i)
    $|v_{k}\rangle$ is horizontally rotated in $\mathcal{E}_{k}$ to $e^{i\theta_{k}}|v_{k}\rangle$ in any given time $\tau$,
  • (ii)
    the curve $\rho_{k;t}=|v_{k;t}\rangle\langle v_{k;t}|$ has the constant Fubini-Study speed $(2\pi\theta_{k}-\theta_{k}^{2})^{1/2}/\tau$.
Condition (i) implies that $H_{t}$ parallel transports the computational space along a loop with holonomy $\Gamma$, and condition (ii) implies that this loop saturates the isoholonomic inequality and hence that the implementation of $\Gamma$ is tight.
The existence of a $H_{k;t}$ satisfying the criteria above can be easily justified geometrically. We begin by studying an evolution of the state $\rho_{k}=|v_{k}\rangle\langle v_{k}|$ of the form
$\rho_{k;t}=e^{-itB_{k}}\rho_{k}e^{itB_{k}},$ (39)
where $B_{k}$ is a Hermitian operator on $\mathcal{E}_{k}$. Choose the eigenbasis of $B_{k}$ such that if $\rho_{k}$ is represented by a Bloch vector $\bm{r}$ and $B_{k}$ by a Rabi vector $\bm{b}$ in $3$-space, then $\bm{r}$ makes an angle $\alpha=\arccos(\theta_{k}/\pi-1)$ with $\bm{b}$; see Figure 2. Adjust the eigenvalues of $B_{k}$ so that $\rho_{k;t}$ becomes periodic with period $\tau$. The curve $\rho_{k;t}$ then has the Fubini-Study speed $(2\pi\theta_{k}-\theta_{k}^{2})^{1/2}/\tau$, and each loop of $\rho_{k;t}$ acquires the geometric phase $\theta_{k}$. However, unless $\theta_{k}=0$ or $\theta_{k}=\pi$, the curve $e^{-itB_{k}}|v_{k}\rangle$ is not horizontal. Therefore, we cannot take $H_{k;t}=B_{k}$.
$\bm{b}$$\bm{r}$$\bm{r}_{t}$$\alpha$
Figure 2: The Bloch vector $\bm{r}$ of $\rho_{k}$ forms an angle $\alpha=\arccos(\theta_{k}/\pi-1)$ with the Rabi vector $\bm{b}$ of the operator $B_{k}$. The Bloch vector $\bm{r}_{t}$ of $\rho_{k;t}$ rotates around $\bm{b}$ and forms a periodic curve with period $\tau$. The angle $\alpha$ is such that each loop of $\rho_{k;t}$ acquires the geometric phase $\theta_{k}$.
To complete the proof of the existence of $H_{k;t}$, we consider the rotating frame picture obtained by applying the transformation $R_{t}=e^{itB_{k}}$ to the system. In this frame, the parallel transport condition, as given by Eq. (16), takes the form:
$\langle v_{k;t}^{\textsc{rf}}|H_{k;t}^{\textsc{rf}}+B_{k}|v_{k;t}^{\textsc{rf}}\rangle=0.$ (40)
We select $H_{k;t}^{\textsc{rf}}$ to be time-independent, $H_{k;t}^{\textsc{rf}}=H_{k}^{\textsc{rf}}$, and choose it such that $|v_{k}^{\textsc{rf}}\rangle=|v_{k}\rangle$ is an eigenvector with eigenvalue $-\langle v_{k}|B_{k}|v_{k}\rangle$. This ensures that the parallel transport condition (40) is satisfied. Also, the evolution equation (39) holds, as $\rho_{k}^{\textsc{rf}}=\rho_{k}$ remains stationary in the rotating frame picture. For more details and a slightly different presentation, see Ref. [22].
$\bm{b}$$\bm{\Omega}^{\textsc{rf}}$$\bm{r}^{\textsc{rf}}$$\bm{b}$$\bm{r}$$\bm{r}_{t}$$\bm{\Omega}$$\bm{\Omega}_{t}$
Figure 3: The left figure shows the Bloch vector $\bm{r}^{\textsc{rf}}$ of $\rho_{k}$ and the Rabi vector $\bm{\Omega}^{\textsc{rf}}$ of $H_{k}^{\textsc{rf}}$ in the rotating frame. Since $\rho_{k}$ is an eigenstate of $H_{k}^{\textsc{rf}}$, $\bm{r}^{\textsc{rf}}$ is parallel to $\bm{\Omega}^{\textsc{rf}}$ and stationary. In the laboratory frame, the Bloch vector $\bm{r}_{t}$ of $\rho_{k;t}$ and the Rabi vector $\bm{\Omega}_{t}$ of $H_{k;t}$ vary with time, as shown in the right figure. The vectors rotate around the axis spanned by the Rabi vector $\bm{b}$ of $B_{k}$ with the same angular speed. That $\rho_{k;t}$ is parallel transported is reflected in the fact that $\bm{r}_{t}$ and $\bm{\Omega}_{t}$ remain perpendicular.
In the laboratory frame,
$H_{k;t}=e^{-itB_{k}}H_{k}e^{itB_{k}},\quad H_{k}=H_{k}^{\textsc{rf}}+B_{k}.$ (41)
Figure 3 shows the Bloch vector of $\rho_{k}$ and the Rabi vector of $H_{k}^{\textsc{rf}}$ in the rotating frame (left panel), and the Bloch vector of $\rho_{k;t}$ and the Rabi vector of $H_{k;t}$ in the laboratory frame (right panel). The Bloch and Rabi vectors are stationary in the rotating frame and rotate with constant angular speed around the axis spanned by $\bm{b}$ in the laboratory frame.
Remark 1.
In the projective case it is sufficient that $\langle v_{k}|H_{k}^{\textsc{rf}}+B_{k}|v_{k}\rangle=\epsilon$ for some real number $\epsilon$. However, $\epsilon$ must be the same for all $k$. In this case, the angle between $\bm{\Omega}_{t}$ and $\bm{r}_{t}$ is $\arccos(\epsilon/|\bm{\Omega}|)$.
V   Summary